A 15 m pole is leaning against a wall. The foot of the pole is 10 m from the wall. What is the angle that the pole makes with the ground?
Answer by nerdybill(7384)

A 15 m pole is leaning against a wall. The foot of the pole is 10 m from the wall. What is the angle that the pole makes with the ground?
.
Apply Trigonometry...
Let x = angle
then
cos x = 10/15
x = cos^(-1)(10/15)
x = cos^(-1)(2/3)
x = 48.2 degrees